# A new approach to constraining properties of AGN host galaxies by combining image and SED decomposition: testing upon the relation

### v1

@@ Abstract
The outshining light from active galactic nuclei (AGNs) poses significant challenges in studying the properties of AGN host galaxies. To address this issue, we propose a novel approach which combines image decomposition and spectral energy distribution (SED) decomposition to constrain properties of AGN host galaxies. Image decomposition allows us to disentangle optical flux into AGN and stellar components, thereby providing additional constraints on the SED models to derive more refined stellar mass. To test the viability of this approach, we obtained a sample of 24 X-ray selected type-I AGNs with redshifts ranging from 0.73 to 2.47. We estimated the stellar masses for our sample and found that our results are generally consistent with earlier estimates based on different methods. Through examining the posterior distribution of stellar masses, we find that our method could derive better constrained results compared to previous SED decomposition methods. With the derived stellar masses, we further studied the relation of our sample, finding a higher intrinsic scatter in the correlation for our entire sample compared to the local quiescent correlation, which could be caused by a few “black hole monsters” in our sample. We propose that based on our method, future works could extend to larger samples of high-redshift AGN host galaxies, thereby enhancing our understanding of their properties.
UTF8gbsn
@@ 1 Introduction
It is widely believed that supermassive black holes (SMBHs) reside in the central regions of most galaxies (e.g., Magorrian et al., 1998; Gebhardt et al., 2001; Kormendy & Ho, 2013). Moreover, tight correlations between BH mass () and properties of the bulge component of the host galaxy (e.g., relation) have been found in the local universe (Ferrarese & Merritt, 2000; Marconi & Hunt, 2003; Häring & Rix, 2004; Bennert et al., 2010; Beifiori et al., 2012; Schramm & Silverman, 2013). When it comes to higher redshifts, separating bulge mass from disk mass becomes difficult, making it harder to study the correlation between BH mass and total stellar mass of the host galaxy (e.g., Ding et al., 2020; Suh et al., 2020; Li et al., 2021a, 2024; Zhuang & Ho, 2023; Tanaka et al., 2024).
@@
In order to explain the origin of this correlation, various theoretical models have been proposed. For example, AGN feedback is invoked in many cosmological simulations, in which a fraction of the AGN energy heats the surrounding gas and regulates the BH accretion and the star formation of the host galaxy (Springel et al., 2005; Di Matteo et al., 2008; Hopkins et al., 2008; DeGraf et al., 2015; Harrison, 2017). Alternatively, introducing an indirect connection in which the BH accretion and star formation share a common gas supply is viable (Cen, 2015; Menci et al., 2016; Anglés-Alcázar et al., 2017). On the other hand, in a cosmic averaging scenario the correlations could be established through mergers based solely on Central Limit Theorem, without involving any close physical connection between the BH and the host galaxy (Peng, 2007; Hirschmann et al., 2010; Jahnke & Macciò, 2011). However, which of these models is the dominant cause of the mass correlation is still in debate.
@@
In order to investigate the potential coevolution, many observational efforts have been put into studying the correlations between BHs and host galaxies for moderate and high redshifts (e.g., Treu et al., 2004, 2007; Jahnke et al., 2009; Schramm & Silverman, 2013; Sun et al., 2015; Park et al., 2015; Suh et al., 2020; Ding et al., 2020; Li et al., 2021a; Zhuang & Ho, 2023; Tanaka et al., 2024). However, obtaining either the BH mass or the properties of the host galaxies is fraught with difficulties. The BH mass can be obtained with dynamical method for nearby galaxies, but it becomes difficult with higher redshifts for the demand of high spatial resolution. Although recently there has been dynamically measured BH mass for a quasar with (Abuter et al., 2024), previous high-redshift BH masses are derived through mainly virial method (e.g., Vestergaard & Peterson, 2006; Schulze et al., 2018) for type-I AGNs. Without the obscuration of the dusty torus according to the unified AGN model (Netzer, 2015), the accretion disk in a type-I AGN glares and outshines the stellar emission of the host galaxy by far in the optical and ultraviolet (UV) bands, bringing difficulty to separating the stellar component from AGN contamination.
@@
However, there have been several methods to estimate the host stellar masses of these unobscured AGNs. For example, Matsuoka et al. (2015) obtained high signal-to-noise ratio (S/N) optical spectra of 191 broad-line quasars with and performed spectral decomposition to derive the velocity dispersion and the mass of the stellar component. UV to far infrared (FIR) broadband SED decomposition (Zou et al., 2019; Suh et al., 2019, 2020; Sun et al., 2024) was employed to derive stellar properties of AGN host galaxies, mainly through adding an AGN template in the SED models. With recent deep imaging surveys, image decomposition techniques have been employed to disentangle host galaxies from core contamination (Peng et al., 2002; Park et al., 2015; Kim et al., 2017; Ding et al., 2020, 2023; Li et al., 2021a, b, 2023, 2024; Zhuang & Ho, 2023; Zhuang et al., 2024). In the image decomposition process, the AGN contribution is modeled as a point spread function (PSF) and subtracted from the original image, leaving the emission from the host galaxies. Ding et al. (2020) estimated stellar masses of 32 X-ray selected AGNs based on stellar fluxes obtained through image decomposition. Setoguchi et al. (2024) used image decomposed SED of the stellar components derived by Li et al. (2021a) to mitigate the contamination in determining BH mass. Inspired by the aforementioned works, here we propose a novel method combining image and SED decomposition to better constrain the stellar mass of host galaxies, where we use the fluxes of the AGNs and the host galaxies derived from image decomposition to refine the corresponding models in the UV-FIR broadband SED fitting.
@@
This paper is structured as follows. Section 2 introduces the sample used in this work. Section 3 describes the Bayesian methodology that we utilized to incorporate additional information from image decomposition to constrain SED models. Section 4 compares our stellar mass estimation with previous works and examines the relation for our sample. Section 5 discusses the probable sources of error in our work and the future prospects of our approach. Finally, Section 6 presents the concluding remarks. Throughout the paper, we assume a flat cosmology (Komatsu et al., 2011) with , and . The luminosities are given in units of .
@@ 2 Sample and data
We initially selected a sample of moderate-luminosity type-I AGNs from Schulze et al. (2018), hereafter S18, in which the near-infrared spectra of the AGN broad line region (BLR) were measured and the black hole masses and spectroscopic redshifts were determined. The sources in our sample are distributed in the CANDELS fields COSMOS, UDS, and GOODS-S, making it easier to search for legacy multi-band data. For the AGNs with black hole mass derived from H or H lines (Trakhtenbrot & Netzer, 2012), we searched for available HST multi-band imaging data and managed to obtain the UV-FIR broadband photometric data for these AGNs from archival catalogs.
We retrieved multi-band HST WFC3/ACS images from the 3D-HST survey (Momcheva et al., 2016). Through crossmatching the S18 catalog and Hubble Source Catalog (Whitmore et al., 2016), we further obtained archival imaging data from MAST Portal. We obtained UV-FIR broadband photometric data for our sample from various archival catalogs, namely GALEX (Martin et al., 2005) for UV, SDSS (Ahumada et al., 2020) and PAN-STARRS (Chambers et al., 2016) for optical, HST (Whitmore et al., 2016) for optical/near-infrared (NIR), UKIRT (Lawrence et al., 2007; Hambly et al., 2008) and VISTA (Cross et al., 2012) for NIR, and WISE (Cutri et al., 2021), Spitzer and Herschel for mid-infrared (MIR) and FIR. We set the matching radius at 1 arcsecond for UV and optical bands. For MIR and FIR bands, given the poorer spatial resolution and pointing precision, we extended the matching radius to 5 arcseconds. Furthermore, we visually inspected the corresponding imaging data and removed the photometric data likely to be contaminated by nearby bright sources.
We employed image decomposition for the matched sources and visually checked the decomposition results. Given the limited imaging quality of the archival data, sources with poorly constrained PSFs tend to yield significantly biased decomposition results and were therefore excluded from further investigation. Our final sample consisted of 24 type-I AGNs, with redshift spanning , black hole mass , Eddington ratio and inferred bolometric luminosity . Figure 1 exhibits the properties of the selected BHs, which aids us in correcting for the selection bias.
@@ 3 Method
In this section we illustrate our new method which combines image decomposition and SED decomposition to help to reduce the degeneracy between AGN bright UV-optical emission and host stellar emission. First the image decomposition procedure is briefly introduced in Section 3.1. Then Section 3.2 describes how we use image decomposition results as extra constraints in the UV-FIR broadband SED decomposition.
@@ 3.1 HST Image decomposition
Following common practice, we simultaneously fitted the two-dimensional flux distribution of the central AGN and the underlying host galaxy, which were modeled as an unresolved point-source component characterized by the PSF and a two-dimensional Sersic profile (e.g., Peng et al., 2002; Park et al., 2015; Kim et al., 2017; Ding et al., 2020; Li et al., 2021a). Although a single Sersic profile may cause uneven residual to be reported when substructures such as spiral arms are present, it acts as an adequate first-order approximation to estimate the total flux of the stellar component, which has been reported in the literature (Matsuoka et al., 2015; Sun et al., 2015; Suh et al., 2020; Ding et al., 2020; Li et al., 2021a).
We used a state-of-the-art image modeling tool Galight (Ding et al., 2021) to perform the AGN-host galaxy image decomposition. Galight is a Python-based open-source package designed for two-dimensional model fitting of especially large samples of extragalactic sources. It makes use of the image modeling capabilities of Lenstronomy (Birrer & Amara, 2018), while adding automatic features such as searching for PSF-stars in the field of view, estimating the noise map of the data and identifying objects to set the initial model and associated parameters.
It has been emphasized that PSF is the dominant source of error in the procedure of flux decomposition (Ding et al., 2020; Tanaka et al., 2024). To minimize the potential effect of PSF-mismatching, we adopted a weighted method from Ding et al. (2020). With the aid of Galight’s automatic PSF searching function, we found isolated and unsaturated point sources within the field of view of each target. We selected the point sources with the top 5 sharpest profiles (i.e. lowest full width at half maximum) as PSFs to be used in the image decomposing routine. We also included the stacked PSFs in the fitting. Using multiple PSFs, we can estimate the uncertainty of the derived fluxes. For a source in a certain band, the weight of each PSF is calculated as
@@
where is the quantity of PSFs used for the image of this band. For those with less than 4 available PSFs, we fix the RMSD as of the mean value as an estimation. Figure 2 shows the decomposition results for ECDFS 391 in 3 bands. The decomposed fluxes of host galaxy and AGN component are shown in Tables 2 and 3 in the appendix.
@@ 3.2 SED fitting
where and ( and ) represent the observational flux and associated uncertainty of galaxy (AGN) component in each of the bands derived from AGN-host image decomposition. This has been implemented in the new version of BayeSED code (Han & Han, 2014, 2018; Han et al., 2023). Compared to the normally defined likelihood in Equation 5, our likelihood additionally constrains the galaxy (AGN) model with the flux obtained through image decomposition.
For the SED modeling of the stellar population, we employed the Bruzual & Charlot (2003) simple stellar population (SSP) model with a Chabrier (2003) stellar initial mass function (IMF), the double power law form of star formation history (SFH), and the Calzetti et al. (2000) dust attenuation law. As in Han et al. (2023), we additionally employ a linear SFH-to-metallicity mapping model to describe the chemical evolution history of the galaxy, which has been found to increase the Bayesian evidence of the SED model. Besides, a cold dust component, which is modelled as a gray body, was added with the energy balance assumption in which the total energy of stellar emission absorbed by cold dust is equal to that reemitted in the IR. Finally, the SED of AGN is modelled as the combination of a clumpy torus model and an empirical SED of type-I quasar where an SMC-like dust attenuation law has been applied.
Figure 3 shows two typical SED fitting results, illustrating the circumstances where the stellar or AGN component dominates respectively.
@@ 4.1 Stellar mass estimation
Here we compare our stellar mass estimates with those from previous studies which employed different techniques on the same sample. Our sample consists of 24 AGNs, including 6 that also appear in the sample studied by Zou et al. (2019), 4 in the sample studied by Suh et al. (2020), and 6 in the sample studied by Ding et al. (2020) (hereafter Z19, S20 and D20). These overlapping sources are illustrated in Figure 4, with the -axis representing our estimates of stellar mass and the -axis denoting those from the corresponding comparison samples. The derived stellar masses of our sample are presented in Table 1.
@@
Z19 and S20 used similar methods to address the degeneracy between AGN and stellar components in the UV-optical bands. By assuming energy conservation for galaxies, they constrained stellar emission with FIR photometry. When compared with Z19, who employed the CIGALE (Noll et al., 2009; Boquien et al., 2018) code in the SED fitting, the median absolute difference between our results and theirs is 0.25 dex. We derived consistent stellar masses for 4 sources with . While for the other 2 sources which have reported by Z19, we derived values lower than 10. For 3 out of 4 overlapping sources, our results are inconsistent with S20, who utilized a custom SED fitting code. We note that our estimates of stellar mass tend to be lower compared to previous SED decomposition results, which could be caused by the additional constraints of image decomposition fluxes that we introduced in our method.
@@
Our estimation is basically consistent with D20, who fitted the decomposed SED with stellar templates to obtain the stellar masses, with median difference of 0.39 dex. However when comparing the points with identical -value in Figure 4, we find that estimates of D20 are basically lower than those given by Z19 and S20. This implies that optical image decomposition and multi-band fluxes may provide inconsistent constraints. Our method combines the advantages of these two methods, thereby the parameter space could be better constrained, which will be discussed in Section 5.2.
Name | RA | DEC | |||
(deg) | (deg) | ||||
(1) | (2) | (3) | (4) | (5) | (6) |
CID 108 | 150.058700 | 2.477378 | 1.257450 | ||
CID 110 | 150.062150 | 2.455014 | 0.728313 | ||
CID 128 | 150.230820 | 2.578165 | 1.326690 | ||
CID 192 | 149.663600 | 2.085205 | 1.329610 | ||
CID 1930 | 150.042450 | 2.629209 | 1.408850 | ||
CID 216 | 149.791790 | 1.872890 | 1.369750 | ||
CID 3570 | 149.641100 | 2.107600 | 1.175990 | ||
CID 40 | 150.199770 | 2.190852 | 1.534250 | ||
CID 597 | 150.526170 | 2.244970 | 1.565880 | ||
CID 607 | 150.609710 | 2.323110 | 1.272450 | ||
CID 644 | 150.511570 | 2.409609 | 1.540710 | ||
CID 71 | 150.123690 | 2.358250 | 0.728576 | ||
CID 72 | 150.091550 | 2.399045 | 2.472040 | ||
CID 79 | 150.173980 | 2.402966 | 0.980806 | ||
CID 87 | 150.133060 | 2.303255 | 1.602540 | ||
ECDFS 379 | 53.112522 | 27.684723 | 0.740331 | ||
ECDFS 391 | 53.124916 | 27.758305 | 1.221270 | ||
LID 1273 | 150.056470 | 1.627495 | 1.564150 | ||
LID 1820 | 149.703580 | 2.578080 | 1.445400 | ||
LID 360 | 150.125110 | 2.861700 | 1.649980 | ||
SXDS 0328 | 34.271282 | 5.274789 | 0.807619 | ||
SXDS 0491 | 34.383537 | 5.225608 | 1.342810 | ||
SXDS 0610 | 34.469402 | 5.259953 | 0.933160 | ||
SXDS 0735 | 34.558102 | 4.878100 | 1.345280 |
@@
Note. — Column (1): object name. Columns (2) and (3): right ascension and declination of the object. Columns (4) and (5): spectroscopic redshift and black hole virial mass given by Schulze et al. (2018). Column (6): median stellar mass derived from Bayesian inference, with lower and upper limits indicating 16th and 84th percentiles. A machine readable version of this table is accessible through this Google drive link.
@@
Based on the derived stellar masses, we further investigated the relation for our sample. Due to the fact that the AGN luminosity is likely to depend on black hole mass on a large extent, selection effects are different for our high-redshift sample and the local anchor (Lauer et al., 2007; Schulze & Wisotzki, 2011), which means that in the high-redshift sample the AGNs with higher black hole masses are more likely to be selected, biasing the relation. For this reason, we firstly employed Monte Carlo Markov Chain (MCMC) simulation to quantify the bias caused by sample selection as Li et al. (2021a) did. Practically, we created a mock catalog of AGNs based on black hole mass function (BHMF) and Eddington Ratio distribution function (ERDF) which were given in Schulze et al. (2015), then performed a luminosity cut to mimic our bolometric luminosity limited sample, as clarified in Figure 1.

### v2

@@ Abstract
The outshining light from active galactic nuclei (AGNs) poses significant challenges in studying the properties of AGN host galaxies. To address this issue, we propose a novel approach which combines image decomposition and spectral energy distribution (SED) decomposition to constrain properties of AGN host galaxies. Image decomposition allows us to disentangle optical flux into AGN and stellar components, thereby providing additional constraints on the SED models to derive more refined stellar mass. To test the viability of this approach, we obtained a sample of 24 X-ray selected type I AGNs with redshifts ranging from 0.73 to 2.47. We estimated the stellar masses for our sample and found that our results are generally consistent with earlier estimates based on different methods. Through examining the posterior distribution of stellar masses, we find that our method could derive better constrained results compared to previous SED decomposition methods. With the derived stellar masses, we further studied the relation of our sample, finding a higher intrinsic scatter in the correlation for our entire sample compared to the local quiescent correlation, which could be caused by a few āblack hole monstersā in our sample. We propose that based on our method, future works could extend to larger samples of high-redshift AGN host galaxies, thereby enhancing our understanding of their properties.
@@ 1 Introduction
It is widely believed that supermassive black holes (SMBHs) reside in the central regions of most galaxies (e.g., Magorrian et al., 1998; Gebhardt et al., 2001; Kormendy & Ho, 2013). Moreover, tight correlations between BH mass () and the properties of the bulge component of the host galaxy (e.g., relation) have been found in the local universe (Ferrarese & Merritt, 2000; Marconi & Hunt, 2003; Häring & Rix, 2004; Bennert et al., 2010; Beifiori et al., 2012; Schramm & Silverman, 2013). When it comes to higher redshifts, separating bulge mass from disk mass becomes difficult, making it harder to study the correlation between BH mass and total stellar mass of the host galaxy (e.g., Ding et al., 2020; Suh et al., 2020; Li et al., 2021a, 2024; Zhuang & Ho, 2023; Tanaka et al., 2024).
@@
In order to explain the origin of this correlation, various theoretical models have been proposed. For example, AGN feedback is invoked in many cosmological simulations, in which a fraction of AGN energy heats the surrounding gas and regulates the BH accretion and the star formation of the host galaxy (Springel et al., 2005; Di Matteo et al., 2008; Hopkins et al., 2008; DeGraf et al., 2015; Harrison, 2017). Alternatively, introducing an indirect connection in which BH accretion and star formation share a common gas supply is viable (Cen, 2015; Menci et al., 2016; Anglés-AlcÔzar et al., 2017). On the other hand, in a cosmic averaging scenario the correlations could be established through mergers based solely on Central Limit Theorem, without involving any close physical connection between the BH and the host galaxy (Peng, 2007; Hirschmann et al., 2010; Jahnke & Macciò, 2011). However, which of these models is the dominant cause of the mass correlation is still in debate.
@@
In order to investigate the potential coevolution, many observational efforts have been put into studying the correlations between BHs and host galaxies for moderate and high redshifts (e.g., Treu etĀ al., 2004, 2007; Jahnke etĀ al., 2009; Schramm & Silverman, 2013; Sun etĀ al., 2015; Park etĀ al., 2015; Suh etĀ al., 2020; Ding etĀ al., 2020; Li etĀ al., 2021a; Zhuang & Ho, 2023; Tanaka etĀ al., 2024). However, obtaining either the BH mass or the properties of the host galaxies is fraught with difficulties. The BH mass can be obtained with the dynamical method for nearby galaxies, but it becomes difficult with higher redshifts because of the demand of high spatial resolution. Although recently BH mass has been dynamically measured for a quasar with (Abuter etĀ al., 2024), previous high redshift BH masses are derived mainly by the virial method (e.g., Vestergaard & Peterson, 2006; Schulze etĀ al., 2018) for type I AGNs. Without the obscuration of the dusty torus according to the unified AGN model (Netzer, 2015), the accretion disk in a type I AGN glares and outshines the stellar emission of the host galaxy by far in the optical and ultraviolet (UV) bands, bringing difficulty to separating the stellar component from AGN contamination.
@@
However, there have been several methods to estimate the host stellar masses of these unobscured AGNs. For example, Matsuoka etĀ al. (2015) obtained high signal-to-noise ratio (S/N) optical spectra of 191 broad-line quasars with and performed spectral decomposition to derive the velocity dispersion and the mass of the stellar component. UV-to-FIR (far-infrared) broadband SED decomposition (Zou etĀ al., 2019; Suh etĀ al., 2019, 2020; Sun etĀ al., 2024) was employed to derive stellar properties of AGN host galaxies, mainly by adding an AGN template in the SED models. With recent deep imaging surveys, image decomposition techniques have been employed to disentangle host galaxies from core contamination (Peng etĀ al., 2002; Park etĀ al., 2015; Kim etĀ al., 2017; Ding etĀ al., 2020, 2023; Li etĀ al., 2021a, b, 2023, 2024; Zhuang & Ho, 2023; Zhuang etĀ al., 2024). In the image decomposition process, the AGN contribution is modeled as a point spread function (PSF) and is subtracted from the original image, leaving the emission from the host galaxies. Ding etĀ al. (2020) estimated stellar masses of 32 X-ray-selected AGNs based on stellar fluxes obtained through image decomposition. Setoguchi etĀ al. (2024) used image decomposed SED of the stellar components derived by Li etĀ al. (2021a) to mitigate contamination in determining the BH mass. Inspired by the aforementioned works, here we propose a novel method combining image and SED decomposition to better constrain the stellar mass of host galaxies, where we use the fluxes of the AGNs and the host galaxies derived from image decomposition to refine the corresponding models in the UV-FIR broadband SED fitting.
@@
This paper is structured as follows. Section 2 introduces the sample used in this work. Section 3 describes the Bayesian methodology that we utilized to incorporate additional information from image decomposition to constrain SED models. Section 4 compares our stellar mass estimation with previous work and examines the relation for our sample. Section 5 discusses the probable sources of errors in our work and the future prospects of our approach. Finally, Section 6 presents the concluding remarks. Throughout the paper, we assume a flat cosmology (Komatsu etĀ al., 2011) with , and . The luminosities are given in units of and the masses are given in units of .
@@ 2 Sample and data
We initially selected a sample of moderate-luminosity type I AGNs from Schulze etĀ al. (2018), hereafter S18, in which the near-infrared spectra of the AGN broad line region (BLR) were measured and the black hole masses and spectroscopic redshifts were determined. The sources in our sample are distributed in the CANDELS COSMOS, UDS, and GOODS-S fields, which makes it easier to search for legacy multiband data. For AGNs with black hole mass derived from H or H lines (Trakhtenbrot & Netzer, 2012), we searched for available HST multi-band imaging data and managed to obtain the UV-FIR broadband photometric data for these AGNs from archival catalogs.
We retrieved multi-band HST WFC3/ACS images from the 3D-HST survey (Momcheva etĀ al., 2016). Through crossmatching the S18 catalog and Hubble Source Catalog (Whitmore etĀ al., 2016), we further obtained archival imaging data from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute111The specific observations analyzed can be accessed via https://doi.org/10.17909/5yh3-w824 (catalog doi: 10.17909/5yh3-w824).. We obtained UV-FIR broadband photometric data for our sample from various archival catalogs, namely GALEX (Martin etĀ al., 2005) for UV, SDSS (Ahumada etĀ al., 2020) and PAN-STARRS (Chambers etĀ al., 2016) for optical, HST (Whitmore etĀ al., 2016) for optical/near-infrared (NIR), UKIRT (Lawrence etĀ al., 2007; Hambly etĀ al., 2008) and VISTA (Cross etĀ al., 2012) for NIR, and WISE (Cutri etĀ al., 2021), Spitzer (Flocāh etĀ al., 2009; Sanders etĀ al., 2007) and Herschel (Griffin etĀ al., 2010; Pilbratt etĀ al., 2010) for mid-infrared (MIR) and FIR. We set the matching radius at 1 arcsecond for UV and optical bands. For MIR and FIR bands, given the poorer spatial resolution and pointing precision, we extended the matching radius to 5 arcseconds. Furthermore, we visually inspected the corresponding imaging data and removed the photometric data likely to be contaminated by nearby bright sources.
We employed image decomposition for the matched sources and visually checked the decomposition results. Given the limited image quality of the archival data, sources with poorly constrained PSFs tend to yield significantly biased decomposition results and were therefore excluded from further investigation. Our final sample consisted of 24 type I AGNs, with redshift spanning , black hole mass , Eddington ratio and inferred bolometric luminosity . Figure 1 shows the properties of the selected BHs, which helps us correct for the selection bias.
@@ 3 Method
In this section, we illustrate our new method which combines image decomposition and SED decomposition to help to reduce the degeneracy between AGN bright UV-optical emission and host stellar emission. First, the image decomposition procedure is briefly introduced in Section 3.1. Then Section 3.2 describes how we use image decomposition results as extra constraints in the UV-FIR broadband SED decomposition.
@@ 3.1 HST Image decomposition
Following common practice, we simultaneously fitted the two-dimensional flux distribution of the central AGN and the underlying host galaxy, which were modeled as an unresolved point-source component characterized by the PSF and a two-dimensional Sersic profile (e.g., Peng etĀ al., 2002; Park etĀ al., 2015; Kim etĀ al., 2017; Ding etĀ al., 2020; Li etĀ al., 2021a). Although a single Sersic profile may cause uneven residuals to be reported when substructures such as spiral arms are present, it acts as an adequate first-order approximation to estimate the total flux of the stellar component, which has been reported in the literature (Matsuoka etĀ al., 2015; Sun etĀ al., 2015; Suh etĀ al., 2020; Ding etĀ al., 2020; Li etĀ al., 2021a).
We used a state-of-the-art image modeling tool Galight (Ding etĀ al., 2021) to perform the AGN-host galaxy image decomposition. Galight is a Python-based open-source package designed for two-dimensional model fitting of especially large samples of extragalactic sources. It makes use of the image modeling capabilities of Lenstronomy (Birrer & Amara, 2018), while adding automatic features such as searching for PSF-stars in the field of view, estimating the noise map of the data and identifying objects to set the initial model and associated parameters.
It has been emphasized that PSF is the dominant source of error in the flux decomposition procedure (Ding etĀ al., 2020; Tanaka etĀ al., 2024). To minimize the potential effect of PSF mismatch, we adopted a weighted method from Ding etĀ al. (2020). With the aid of Galightās automatic PSF searching function, we found isolated and unsaturated point sources within the field of view of each target222In the processing procedure for a target source, only the PSFs within the same band, near the target, are involved.. After visual inspection the profiles with poor SNR are excluded, then we selected the point sources with the top 5 sharpest profiles (i.e. the lowest full width at half maximum) as PSFs to be used in the image decomposing routine. We also included stacked PSFs in the fitting. Using multiple PSFs, we can estimate the uncertainty of the derived fluxes. For a source in a certain band, the weight of each PSF is calculated as
@@
where is the quantity of PSFs used for the image of this band. For those with less than four available PSFs, we fix the RMSD as of the mean value as an estimation. Figure 2 shows the decomposition results for ECDFS 391 in three bands. The decomposed fluxes of the host galaxy and the AGN component are shown in Tables 2 and 3 in the appendix.
@@ 3.2 SED fitting
where and ( and ) represent the observational flux and associated uncertainty of galaxy (AGN) component in each of the bands derived from AGN-host image decomposition. This has been implemented in the new version of the BayeSED code (Han & Han, 2014, 2018; Han etĀ al., 2023). Compared to the normally defined likelihood in Equation 5, our likelihood additionally constrains the galaxy (AGN) model with the flux obtained through image decomposition.
For the SED modeling of the stellar population, we employed the Bruzual & Charlot (2003) simple stellar population (SSP) model with a Chabrier (2003) stellar initial mass function (IMF), the double power law form of star formation history (SFH), and the Calzetti etĀ al. (2000) dust attenuation law. As in Han etĀ al. (2023), we additionally employ a linear SFH-to-metallicity mapping model to describe the chemical evolution history of the galaxy, which has been found to increase the Bayesian evidence of the SED model. In addition, a cold dust component, which is modeled as a gray body, was added with the energy balance assumption, in which the total energy of stellar emission absorbed by cold dust is equal to that reemitted in the IR. Finally, the SED of AGN is modelled as the combination of a clumpy torus model and an empirical SED of type I quasar where an SMC-like dust attenuation law has been applied.
Figure 3 presents 4 representative SED fitting results, demonstrating various scenarios for decomposing AGNs and host galaxies. A user-friendly documentation 333https://github.com/hanyk/BayeSED3/blob/main/observation/agn_host_decomp/demo.ipynb is available on the homepage of BayeSED, where readers can access the multi-component photometric data to reproduce our SED fitting results.
@@ 4.1 Stellar mass estimation
Here we compare our stellar mass estimates with those from previous studies which employed different techniques on the same sample. Our sample consists of 24 AGNs, including 6 that also appear in the sample studied by Zou etĀ al. (2019), 4 in the sample studied by Suh etĀ al. (2020), and 6 in the sample studied by Ding etĀ al. (2020) (hereafter Z19, S20, and D20). These overlapping sources are illustrated in Figure 4, with the -axis representing our estimates of stellar mass and the -axis denoting those from the corresponding comparison samples. The derived stellar masses of our sample are presented in Table 1.
@@
Z19 and S20 used similar methods to address the degeneracy between AGN and stellar components in the UV-optical bands. By assuming energy conservation for galaxies, they constrained stellar emission with FIR photometry. Compared with Z19, who employed the CIGALE (Noll etĀ al., 2009; Boquien etĀ al., 2018) code in the SED fitting, the median absolute difference between our results and theirs is 0.27 dex. We derived consistent stellar masses for 4 sources with . However, for the other 2 sources that have reported by Z19, we derived slightly lower values. For three of four overlapping sources, our results are inconsistent with S20, who utilized a custom SED fitting code. We note that our estimates of stellar mass tend to be lower compared to previous SED decomposition results, which could be caused by the additional constraints of image decomposition fluxes that we introduced in our method.
@@
Our estimation is basically consistent with D20, who fitted the decomposed SED with stellar templates to obtain the stellar masses, with a median difference of 0.21 dex. However, when comparing the points with identical -value in Figure 4, we find that the estimates of D20 are basically lower than those given by Z19 and S20. This implies that optical image decomposition and multiband fluxes may provide inconsistent constraints. Our method combines the advantages of these two methods. Therefore, the parameter space could be better constrained, which will be discussed in Section 5.2.
| Name | RA | DEC | |||||
| (deg) | (deg) | ||||||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
| CID 108 | 150.058685 | 2.477427 | 1.257446 | 45.61 | |||
| CID 110 | 150.062134 | 2.455013 | 0.728313 | 44.91 | |||
| CID 128 | 150.230789 | 2.578183 | 1.408853 | 45.96 | |||
| CID 192 | 149.663589 | 2.085209 | 1.220132 | 45.44 | |||
| CID 1930 | 150.042450 | 2.629209 | 1.566184 | 45.74 | |||
| CID 216 | 149.791779 | 1.872894 | 1.565879 | 45.04 | |||
| CID 3570 | 149.641083 | 2.107653 | 1.242593 | 45.34 | |||
| CID | 150.199753 | 2.190866 | 1.534250 | 45.69 | |||
| CID 597 | 150.526184 | 2.244959 | 1.272449 | 45.50 | |||
| CID 607 | 150.609695 | 2.323107 | 1.294432 | 45.83 | |||
| CID 644 | 150.511566 | 2.409609 | 0.984967 | 45.07 | |||
| CID 71 | 150.123672 | 2.358294 | 0.728576 | 44.91 | |||
| CID | 150.091537 | 2.399079 | 2.472042 | 45.96 | |||
| CID 79 | 150.173965 | 2.402994 | 0.980806 | 45.07 | |||
| CID 87 | 150.133041 | 2.303283 | 1.602544 | 45.36 | |||
| ECDFS 379 | 53.112499 | 27.684805 | 0.740331 | 45.43 | |||
| ECDFS 391 | 53.124916 | 27.758333 | 1.221274 | 45.73 | |||
| LID 1273 | 150.056473 | 1.627495 | 1.620638 | 45.77 | |||
| LID 1820 | 149.703598 | 2.578055 | 1.536443 | 45.77 | |||
| LID 360 | 150.125092 | 2.861743 | 1.582554 | 46.03 | |||
| SXDS 0328 | 34.271252 | 5.274861 | 0.807619 | 44.99 | |||
| SXDS 0491 | 34.383335 | 5.225667 | 1.342809 | 45.67 | |||
| SXDS 0610 | 34.469166 | 5.259972 | 0.933160 | 45.09 | |||
| SXDS 0735 | 34.557919 | 4.878194 | 1.442290 | 45.85 |
@@
Note. ā : Sources with BH properties derived from H line. Column (1): object name. Columns (2) and (3): right ascension and declination of the object. Columns (4) (5) (6) and (7): spectroscopic redshift, bolometric luminosity, eddington ratio and BH mass given by (Schulze etĀ al., 2018). Column (6): median stellar mass derived from Bayesian inference, with lower and upper limits indicating 16th and 84th percentiles. A machine readable version of this table is accessible through this google drive link.
@@
Based on the derived stellar masses, we further investigated the relation for our sample. Due to the fact that the AGN luminosity is likely to depend on black hole mass to a large extent, the selection effects are different for our high redshift sample and the local anchor (Lauer etĀ al., 2007; Schulze & Wisotzki, 2011), which means that in the high-redshift sample the AGNs with higher black hole masses are more likely to be selected, biasing the relation. For this reason, we first used Monte Carlo Markov Chain (MCMC) simulation to quantify the bias caused by sample selection as Li etĀ al. (2021a) did. Practically, we created a mock catalog of AGNs based on black hole mass function (BHMF) and Eddington Ratio distribution function (ERDF) which were given in Schulze etĀ al. (2015), then performed a luminosity cut to mimic our bolometric luminosity limited sample, as clarified in Figure 1.